OpenStudy (anonymous):

What is the derivative of sin(1/t)?

terenzreignz (terenzreignz):

Chain rule comes to mind?

OpenStudy (anonymous):

Yes, however I'm confused regarding the sin

terenzreignz (terenzreignz):

Refresher: \[\Large \frac{d}{dt}f[g(t)]= f'[g(t)]g'(t)\]

OpenStudy (anonymous):

Yes I understand that. I'm wondering if I would use the quotient rule for (1/t)

terenzreignz (terenzreignz):

You can, or you can also just use the power rule and recall that \[\Large \frac1t= t^{-1}\]

OpenStudy (anonymous):

Ah alright. That's what I always miss. I think I can get the rest. Thanks :D

terenzreignz (terenzreignz):

Good. Post your answers if you can :D

OpenStudy (anonymous):

Will do.

OpenStudy (anonymous):

So the derivative is cos(1/t)(-1t^-2) right?

terenzreignz (terenzreignz):

That is correct. Just try to write that more neatly (maybe do away with the negative exponent and use a fraction instead) but for the most part, well done ^_^

terenzreignz (terenzreignz):

(in my humble opinion) this looks better: \[\Large -\frac{\cos\left(\frac1t\right)}{t^2}\]

OpenStudy (anonymous):

Yes, that is how my text book would state it. I will get used to doing that :D

OpenStudy (anonymous):

Also, I'm doing this for a velocity problem, and a follow up question is when would the velocity be it's greatest. Can you help me out with understanding that part?

OpenStudy (anonymous):

v(t) = sin(1/t) was the equation for the velocity and the answer I just gave you was for acceleration.

OpenStudy (anonymous):

Here's the full question

OpenStudy (anonymous):

A particle moves along the x-axis so that at any time t, 0.1<0.3, its velocity is given by v(t) = sin(1/t). How would I find when the velocity is at the greatest?